# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v2

@@ B.2 Proof of Theorem 2
Recall the definition $\nabla F({\bm{w}}_{t}^{(g)}),\widehat{\nabla}F({\bm{w}}_{t}^{(g)})$ in Lemma and eq (91), the gradient approximation error is
@@
where step (a) comes from Young’s inequality with $\zeta>0$ to be chosen and step (b) follows from the step size condition $\gamma\leq 1/L_{g}$.
@@ B.3 Proof of Corollary 1
B.4 Proof of The Local Model Convergence Rate in Theorem 2
@@ C.1 Additional Experiment Details
Selection of $\lambda$. For the logistic regression on the synthetic dataset, we selected three values for $\lambda$: small $\lambda=0.02$, medium $\lambda=0.1$, and large $\lambda=0.5$. For logistic regression on the MNIST dataset, we used small $\lambda=0.5$, medium $\lambda=1.5$, and large $\lambda=2.5$. For CNN on the MNIST dataset, we used small $\lambda=0.2$, medium $\lambda=0.5$, and large $\lambda=2.5$. Such a choice is to make sure that each line in the figure is clearly separable. Additionally, we implemented an adaptive choice of $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$ following the formula outlined in Theorem 1, with $\rho$ set to 2. The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance.
@@
Selection of Step Size. For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=\frac{\lambda+L}{\lambda L}$ and the local step size $\eta=\frac{1}{L+\lambda}$. The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$. The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$. For the MNIST dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search. For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$. LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
@@
Selection of Hyperparameters. For the MNIST dataset, we implemented logistic regression using the SGD solver with 5 epochs, a batch size of 32, and 20 total runs. For the CNN model on the MNIST dataset, the same hyperparameter setup was used, except the batch size was reduced to 16 to ensure stable convergence.
@@
Optimizer. For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD). For the MNIST dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
@@ C.2 Additional Experiment Results
In Figure 3, Statistical error of the FedProx algorithm with an adaptive choice of $\lambda$ (from Theorem 1) under varying levels of statistical heterogeneity. With the adaptive choice of $\lambda$ as defined in Theorem 1, FedProx consistently achieves the lowest statistical error across most scenarios, outperforming both GlobalTrain and LocalTrain baselines. This is particularly evident under moderate and high levels of statistical heterogeneity. However, as $R$ approaches very small values, the global baseline performs slightly better, which aligns with the slower rate of FedProx with small $R$ described in our theoretical results. Overall, these results validate our choice of adaptive $\lambda$ and confirm the statistical rate provided by our analysis.

### v3

@@ B.3 Proof of Theorem 2
Recall the definition $\nabla F({\bm{w}}_{t}^{(g)}),\widehat{\nabla}F({\bm{w}}_{t}^{(g)})$ in Lemma 3 and eq (92), the gradient approximation error is
@@
where the step (a) comes from the Young’s inequality with $\zeta>0$ to be chosen and the step (b) follows from the step size condition $\gamma\leq 1/L_{g}$.
@@ B.4 Proof of Corollary 1
B.5 Proof of The Local Model Convergence Rate in Theorem 2
@@ C.1 Additional Experiment Details
Real Dataset. The MNIST dataset consists of 70,000 grayscale images of handwritten digits, each 28x28 pixels, classified into 10 classes (digits 0-9). The EMNIST dataset extends MNIST, including handwritten letters and digits. The Balanced split contains 131,600 28x28 grayscale images across 47 balanced classes (digits and uppercase/lowercase letters). And the CIFAR-10 dataset contains 60,000 color images (32x32 pixels, 3 RGB channels), categorized into 10 classes representing real-world objects (e.g., airplane, car, and dog).
Algorithms. To fully investigate the effect of personalization in FedProx with the Algorithm B.1, we compare the two extreme algorithms (GlobalTraining and LocalTraining). Furthermore, since there exist other methods solving the objective (1) to do personalized federated learning, we also try to investigate the effect of personalization in other methods, like pFedMe (T Dinh et al., 2020). The algorithm pFedMe has a similar design with a double loop to solve the subproblem in each communication round.
@@
Selection of $\lambda$. For the logistic regression on the synthetic dataset, we selected three values for $\lambda$: small $\lambda=0.02$, medium $\lambda=0.1$, and large $\lambda=0.5$. For logistic regression on the MNIST dataset, we used small $\lambda=0.5$, medium $\lambda=1.5$, and large $\lambda=2.5$. For CNN on the MNIST dataset, we used small $\lambda=0.2$, medium $\lambda=0.5$, and large $\lambda=2.5$. Such a choice is to make sure that each line in the figure is clearly separable. Additionally, we implemented an adaptive choice of $\lambda$ following the formula outlined in Theorem 1, with $\rho$ set to 2. The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance. For other real datasets, the regularization term $\lambda$ is set as small $\lambda=0.1$, medium $\lambda=$ 0.5, and large $\lambda=$ 1.
@@
Selection of Step Size. For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=\frac{\lambda+L}{\lambda L}$ and the local step size $\eta=\frac{1}{L+\lambda}$. The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$. The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$. For the real dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search. For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$. LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
@@
Selection of Hyperparameters and Convolutional Neural Network Model. For the MNIST dataset, we implemented logistic regression using the SGD solver with 5 epochs, a batch size of 32, and 20 total runs, and the same hyperparameter setup was used, except the batch size was reduced to 16 to ensure stable convergence in CNN with two convolution layers. To make our experiment more comprehensive, we also utilized a five-layer CNN to train the MNIST dataset with 10 epochs, a batch size of 16, and 50 total runs. For the EMNIST dataset, we used the same five-layer CNN with the same settings. And for CIFAR10, we implemented a three-layer CNN with 10 epochs, a batch size of 16, and 50 total runs. To study the data heterogeneity, we set the number of classes for each client in the MNIST, EMNIST, and CIFAR10 datasets as $\{2,6,10\},\{30,40,47\},$ and $\{2,6,10\}$, respectively.
@@
Optimizer. For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD). For the real dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
@@ C.2 Additional Experiment Results
In Figure 3, statistical error of the FedProx algorithm with an adaptive choice of $\lambda$ under varying levels of statistical heterogeneity. With the adaptive choice of $\lambda$ as defined in Theorem 1, FedProx consistently achieves the lowest statistical error across most scenarios, outperforming both GlobalTrain and LocalTrain baselines. This is particularly evident under moderate and high levels of statistical heterogeneity. However, as $R$ approaches very small values, the global baseline performs slightly better, which aligns with the slower rate of FedProx with small $R$ described in our theoretical results. Overall, these results validate our choice of adaptive $\lambda$ and confirm the statistical rate provided by our analysis.
@@
Dataset Client Class One-stage FedProx GlobalTrain LocalTrain pFedMe
(small) (medium) (large) (small) (medium) (large)
MNIST CNN2 2 0.3567 0.4721 0.6019 0.6123 0.3019 0.3973 0.5828 0.6172
6 0.7001 0.7833 0.7843 0.8092 0.6231 0.7918 0.8093 0.8312
10 0.8753 0.8893 0.9032 0.9098 0.8194 0.8756 0.8771 0.8749
MNIST CNN5 2 0.8891 0.9000 0.9333 0.9377 0.6536 0.9000 0.9003 0.9380
6 0.8451 0.8941 0.9392 0.9446 0.7333 0.9288 0.9186 0.9306
10 0.9091 0.9102 0.9421 0.9433 0.7633 0.9340 0.9385 0.9440
EMNIST CNN5 30 0.6084 0.6089 0.6011 0.6102 0.5583 0.6133 0.6129 0.6122
40 0.6097 0.6123 0.6357 0.6423 0.5644 0.6012 0.6102 0.6134
47 0.6278 0.6415 0.6672 0.6611 0.6111 0.6345 0.6532 0.6717
CIFAR10 CNN3 2 0.6101 0.6292 0.6340 0.6444 0.5801 0.6056 0.6012 0.6033
6 0.6211 0.6311 0.6712 0.6949 0.5623 0.6163 0.6163 0.6439
10 0.6102 0.6712 0.7302 0.8041 0.6238 0.6744 0.7443 0.7732
[…]
To further validate our conclusions about the effect of personalization on the statistical accuracy of FedProx (Objective 1), we conduct additional experiments on real datasets, including EMNIST and CIFAR-10. For each dataset, we explore three levels of statistical heterogeneity, controlled by varying the number of classes each client has access to, which follows a widely used setup in federated learning literature. The models are trained using different convolutional neural network (CNN) architectures, denoted as CNN2 and CNN5, where the numbers indicate the number of convolutional layers in the network. As the true underlying model is not accessible for these real datasets, we evaluate model performance based on classification accuracy on the test set, as shown in Tables 3 and 4. We also conduct experiments another regularization-based personalization method, pFedMe (T Dinh et al., 2020), to demonstrate that our results could be generalized to other regularization-based personalized federated learning methods as well. For the columns corresponding to One-stage FedProx and pFedMe, we conduct experiments with varying regularization parameter $\lambda$, investigating the role of personalization in the algorithms. Specifically, we experiment with small, medium, and large $\lambda$ values (see Appendix C.1 for detailed parameter settings).
In Table 3, we can see that consistently across datasets, as $\lambda$ decreases (indicating a higher degree of personalization), the accuracy approaches that of LocalTrain, which trains separate models for each client. Conversely, as $\lambda$ increases (indicating lower personalization and more collaboration), the accuracy approaches that of GlobalTrain, which represents purely global training. This interpolation behavior between LocalTrain and GlobalTrain, enabled by FedProx’s personalized objective, aligns well with our theoretical results in Theorem 1 and highlights the adaptive capability of the FedProx method.
Moreover, as dataset homogeneity increases (i.e., clients share more similar data distributions), we observe a general improvement in accuracy across all methods, regardless of the level of personalization. This result demonstrates that while personalization offers significant benefits in highly heterogeneous settings, its necessity diminishes as heterogeneity decreases. Additionally, experiments with different CNN architectures confirm the robustness of these findings across model designs. Although we did not provide a theoretical analysis for pFedMe in this paper, our results suggest that the empirical behavior of pFedMe is consistent with the theoretical insights derived for FedProx. In particular, we observe that pFedMe also exhibits interpolation between local and global learning under varying levels of data heterogeneity.
Dataset Client Class One-stage FedProx (small $\lambda$) One-stage FedProx (large $\lambda$) Dichotomous Strategy
MNIST CNN2 2 0.4721 0.6019 0.2567
MNIST CNN2 6 0.7833 0.7843 0.7001
MNIST CNN2 10 0.8893 0.9032 0.8753
MNIST CNN5 2 0.8891 0.9333 0.7032
MNIST CNN5 6 0.8451 0.8941 0.7941
MNIST CNN5 10 0.9091 0.9102 0.8922
EMNIST CNN5 30 0.6084 0.6089 0.5291
EMNIST CNN5 40 0.6097 0.6123 0.5732
EMNIST CNN5 47 0.6278 0.6415 0.5856
Table 4: Comparison of evaluation accuracy for One-stage FedProx with different values of $\lambda$ and Dichotomous Strategy across various datasets and client classes.
C.3 Comparison of FedProx with A Dichotomous Strategy
A simple baseline in personalized federated learning is the dichotomous strategy, where the dataset on each client is split into two subsets: one is used for global training (GlobalTrain) with pooled client data, while the other is used for purely local training (LocalTrain). After training, the two models are compared on the test set, and the best-performing model is selected for each client. Such a strategy is known to achieve minimax-optimal statistical accuracy under certain conditions; see Chen et al. (2023b) for a detailed discussion.
Even though the convergence rate coincides with such a dichotomous strategy, we may reiterate that the key contribution of our work lies in the fine-grained analysis of the FedProx method—a method that has demonstrated strong empirical performance in federated learning [2]. Through our analysis, we have, for the first time, established the minimax optimality of FedProx, showing that its statistical performance matches the lower bound of the estimation problem. If one seeks to understand why FedProx could be advantageous over the dichotomous strategy in practice, here are some of the reasons:
  • Stability in Heterogeneous Environments: Empirical studies [1] demonstrate that FedProx achieves more stable convergence compared to methods such as FedAvg, especially in settings with significant in a heterogeneous network.
  • Adaptivity and Flexibility: FedProx seamlessly interpolates between the extremes of global training and local training by tuning a single regularization parameter, $\lambda$. This adaptability allows practitioners to dynamically balance the trade-off between bias and variance based on the level of statistical heterogeneity, without the need for fixed, hard decisions inherent in the dichotomous strategy.
To further compare FedProx with the dichotomous strategy, we conduct a simple experiment. In the dichotomous setup, each client’s dataset is partitioned into two subsets: one subset is used for local training (LocalTrain), where each client trains independently, and the other subset is pooled across all clients for global training (GlobalTrain). The accuracy of the two resulting models is evaluated on a test set, and the better-performing model is reported for each client. By contrast, FedProx is implemented using the experimental setup described earlier, leveraging the full dataset. As shown in Table 4, FedProx consistently outperforms the dichotomous strategy across various configurations. This performance gap can be attributed to the inefficiency of the dichotomous approach, which sacrifices valuable data during the sample splitting process, resulting in reduced sample efficiency.
Even in a hypothetical scenario where the full sample is used for both LocalTrain and GlobalTrain in the dichotomous strategy, the results in Table 3 demonstrate that FedProx achieves a continuous interpolation between the two extremes of LocalTrain and GlobalTrain by adjusting its regularization parameter $\lambda$. This aligns with our theoretical findings.